OpenStudy (anonymous):

Y – 64 = m(x – 64) what is this in slope intercept form? I'm having some difficulty solving it.

OpenStudy (anonymous):

Really what I wanted to ask was what to do when you distribute m to x and 64.

hartnn (hartnn):

a(b-c) = ab-ac m(x-64) = mx-64m